Suppose you have a portfolio that is 30% in the risk-free asset and 70% in a stock. The stock has a standard deviation of 0.30 (i.e., 30%). What is the standard deviation of the portfolio?

a) 0.30 (30%)

B) 0.09 (9%)

C) 0.21 (21%)

D) 0
Answer #1

Standard deviation of portfolio is mathematically represented as:

Oportfolio = wig+wo3 + 2wW2P1,20102 Where: W W 04 02 P12 = = = = = Proportion of the portfolio invested in Asset 1 Proportion

STandard deviation of risk free asset and its correlation with any other stock returns' is = 0

0= V(0.30 * 0)2 + (0.70*0.30)2 + (2 * 0.30*0*0.70*0.30 * 0)

o= V(0.70 * 0.30)

0 = 0.21 (Option C)
Answer #2

C) 0.21 (21%)